# Mut4All: Fuzzing Compilers via LLM-Synthesized Mutators Learned from Bug Reports

### v1

@@ 4.2. Evaluated Languages and Compilers
Language Compiler Fuzzer Avenue
Rust rustc/gccrs RustSmith (Sharma et al., 2023) ISSTA-23
Rustlantis (Wang and Jung, 2024) OOPSLA-24
Clozemaster (Gao et al., 2025) ICSE-25
C++ GCC/Clang CSmith (Yang et al., 2011) PLDI-11
YARPGen-2.0 (Livinskii et al., 2023) PLDI-23
GrayC (Even-Mendoza et al., 2023) ISSTA-23
Fuzz4All (Xia et al., 2024) ICSE-24
MetaMut (Ou et al., 2024) ASPLOS-25
For Rust, we adopt rustc [^6] https://www.rust-lang.org/ (the official compiler) and gccrs [^7] https://github.com/Rust-GCC/gccrs/ (the GNU front-end for Rust). For C++, we use GCC [^8] https://gcc.gnu.org/ and Clang, [^9] https://clang.org/ which are the two most widely deployed C++ compilers in practice. We use the latest release versions of all compilers: rustc 1.88, gccrs-a1a56c6, GCC 14.1.0, and Clang 18.1.0. As test oracles, we apply crash, hang, and differential testing for compilers of the same language (Gao et al., 2025; Zhang et al., 2017). Among them, crash and hang are used to detect compiler failures, the former indicates a crash or assertion failure, while the latter indicates an infinite loop or deadlock. Differential testing is used to detect inconsistent behaviors between different compilers, which is a common practice for testing C/C++ compilers that share the same language standard but may have different implementations. For the baseline fuzzers, we use their default configurations from their official repositories. For approaches that require seed programs, we use the same initial seed corpus as Mut4All to ensure a fair comparison (i.e., the enhanced seeds).
@@ 4.2.1. Baselines
To thoroughly evaluate the effectiveness of Mut4All, we compare it against both traditional compiler fuzzers and recent LLM-based approaches. In particular, to investigate how our mutator-driven pipeline compares to direct LLM-based mutated program generation, we include the latest LLM-based fuzzers that modify program source code directly—namely, Clozemaster (Gao et al., 2025) for Rust and Fuzz4All (Xia et al., 2024) for C++. In total, we evaluate three fuzzers for Rust and five for C++, as shown in Table 1.
Traditional Fuzzers.  For Rust, we adopt generation-based fuzzers such as RustSmith (Sharma et al., 2023) and Rustlantis (Wang and Jung, 2024). For C++, we include CSmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023), which follow a similar generation-based paradigm by randomly expanding language grammars and production rules to produce valid programs. Additionally, we include GrayC (Even-Mendoza et al., 2023), a representative mutation-based C++ fuzzer equipped with manually designed, general-purpose mutators.
LLM-based Fuzzers.  We further compare Mut4All with recent LLM-based compiler fuzzers that follow different design philosophies: we let the LLMs design and implement mutators, which are then applied to the seed programs but they let the LLMs directly edit the source code to generate mutated programs. For Rust, Clozemaster (Gao et al., 2025) fine-tunes an LLM to complete masked code segments, enabling direct in-place edits to source programs. For C++, Fuzz4All (Xia et al., 2024) utilizes LLMs to rewrite seed programs guided by language standards and official API references. In contrast, MetaMut (Ou et al., 2024) uses an LLM to systematically synthesize and combine simple transformation rules into new mutators, falling into the same category as Mut4All but with limited mutator diversity.
@@ 4.3. RQ1: Comparison with Existing Fuzzers
To evaluate Mut4All against existing approaches, we run each fuzzer with a 24-hour time budget, which is a common setting in prior work (Even-Mendoza et al., 2023; Xia et al., 2024; Ou et al., 2024; Yang et al., 2024a). Also, each fuzzer is executed independently for five trials. Fuzz4All and MetaMut both use GPT-4o as Mut4All, while Clozemaster employs a fine-tuned model. We execute each fuzzer five times independently.
4.3.1. Unique Crashes
Following MetaMut (Ou et al., 2024), we identify a unique crash based on the top-2 frames of the stack trace, excluding compiler helper functions. We aggregate all unique crashes across the five runs for each fuzzer, as shown in Figure 7.
Mut4All consistently discovers the highest number of unique crashes for both Rust and C++, significantly surpassing all baselines. For Rust (Figure 7(a)), Mut4All uncovers 60 unique crashes, whereas Clozemaster and Rustlantis report only 5 and 0, respectively. Notably, only 6 crashes overlap between Mut4All and Clozemaster, indicating that the vast majority of bugs found by Mut4All are exclusive. For C++, Mut4All detects 35 unique crashes, compared to 27 from MetaMut and 10 from Fuzz4All. The overlap is minimal—only 3 crashes are shared between Mut4All and each of these fuzzers—demonstrating that most of Mut4All’s findings are novel. In contrast, traditional fuzzers such as CSmith, GrayC, and YARPGen report no unique crashes in our experiments, suggesting that these conventional techniques are reaching their saturation point.
(a) Rust
(b) C++
Figure 7. Venn diagram of discovered unique crashes.
@@ 4.3.2. Coverage
Figure 8. Line coverage over 24h of fuzzing.
Figure 8 presents the average coverage trends of all evaluated fuzzers. For Rust (Figure 8(a)), Mut4All achieves the highest line coverage on rustc, outperforming Clozemaster and Rustlantis. For C++ compilers, MetaMut achieves the highest code coverage on both GCC (Figure 8(b)) and Clang (Figure 8(c)). Mut4All ranks second, followed closely by Fuzz4All. This result aligns with expectations, as MetaMut employs a coverage-guided fuzzing strategy that explicitly optimizes for maximizing code paths executed. In contrast, Mut4All focuses on synthesizing and evaluating a diverse set of mutators across seed programs, without using coverage feedback as a direct guidance signal. Despite this, Mut4All still achieves competitive coverage when compared to state-of-the-art fuzzers, particularly in Rust. Mut4All ’s diverse mutators improve coverage without explicit optimization for coverage.
Answer to RQ1: Mut4All consistently outperforms all baselines in finding unique crashes, and achieves the best coverage on Rust and the second-best on C++.
@@ 4.4.1. The Status of Our Reported Bugs
We have synthesized 319 and 403 valid mutators for Rust and C++, respectively. Using these mutators, we conducted a 48-hour fuzzing campaign for each language and reported a total of 96 bugs, as summarized in Table 4.4.1. Specifically, we identified 62 bugs in Rust compilers (44 in rustc and 18 in gccrs) and 34 in C++ compilers (21 in GCC and 13 in Clang). Among these, 54 bugs (56%) have been confirmed by compiler maintainers as previously unknown bugs: 33 in rustc, 5 in gccrs, 6 in GCC, and 10 in Clang. Notably, 8 of the confirmed bugs have already been fixed by the respective compiler developers. In addition, 14 of our reports were classified as duplicates of existing bugs, and 2 rustc reports were marked as Won’t Fix.
@@
Table 2. Status of the reported bugs by Mut4All
@@
Compiler rustc gccrc GCC Clang
@@
Confirmed as New 33 5 6 10 54
Fixed 4 3 0 1 8
Duplicate 9 0 2 3 14
@@ 4.4.2. Bug Examples
We further present two more bugs to show Mut4All’s effectiveness in uncovering subtle, hard-to-construct compiler issues beyond prior fuzzers.
@@
4impl<V, D> Trait<X, X> for A // ‘Trait<V, D>‘ -> ‘Trait<X, X>‘
5where A: Trait<D, V>, B: Trait<V, D>, {}
6impl<V, D> Trait<X, X> for B // ‘Trait<V, D>‘ -> ‘Trait<X, X>‘
7where A: Trait<V, D> {}
8fn impls_trait<T: Trait<V, D>, V, D>() {}
9fn main() { impls_trait::<B, _, _>(); }
Listing 6: The new bug Rust-14XX63 uncovered by Mut4All.
Listing 6 shows a Rust program where valid trait parameters are replaced with an undefined type (i.e., X), causing rustc to enter an infinite loop due to unresolved recursive bounds. While such mutations are unlikely to be crafted manually, Mut4All automatically identifies them, exposing a deep bug in the compiler’s trait resolution logic.
@@
Listing 7: The new bug LLVM-14XX70 uncovered by Mut4All.
Listing 7 presents a C++ program using template and constexpr. A mutator from Mut4All modifies a valid function fooT() into one with an invalid parameter list: fooT(C auto,). Though syntactically incorrect, this triggers a Clang crash, exposing a long-standing bug dating back to version 9.0 (i.e., more than 6 years). Such unintuitive mutations in sensitive constructs are rarely produced by traditional fuzzers like MetaMut, nor by human developers. Mut4All’s systematic exploration of these edge cases demonstrates its ability to uncover deep compiler bugs.
@@
Answer to RQ2: Mut4All discovered 96 bugs across four production compilers, 54 of which have been confirmed as previously unknown, demonstrating its effectiveness in hunting real compiler bugs.
4.5. RQ3: Mutator Synthesis Effectiveness
The mutators designed and implemented in Mut4All have shown their effectiveness in fuzzing compilers. In this RQ, we evaluate the overall generation cost and the common problems in valid/invalid mutators, which can help us understand the effectiveness of Mut4All in mutator synthesis.
4.5.1. Overall Effectiveness and Cost
We sampled 500 historical bug reports for both Rust and C++ compilers, and used them to construct prompts for Mut4All to invent and synthesize mutators. We then performed a statistical analysis of the synthesis outcomes.

### v2

@@ 4.2. Evaluated Languages and Compilers
Language Compiler Fuzzer Avenue Description
Rust rustc/gccrs RustSmith (Sharma et al., 2023) ISSTA-23 Traditional Generation-Based
Rustlantis (Wang and Jung, 2024) OOPSLA-24 Traditional Generation-Based
Clozemaster (Gao et al., 2025) ICSE-25 LLM-Based Directly Generation Fuzzer
C++ GCC/Clang CSmith (Yang et al., 2011) PLDI-11 Traditional Generation-Based
YARPGen-2.0 (Livinskii et al., 2023) PLDI-23 Traditional Generation-Based
GrayC (Even-Mendoza et al., 2023) ISSTA-23 Traditional Mutation-Based
Fuzz4All (Xia et al., 2024) ICSE-24 LLM-Based Directly Generation Fuzzer
MetaMut (Ou et al., 2024) ASPLOS-24 LLM-Based Mutation-Based Fuzzer
TyMut (Wang et al., 2025b) OOPSLA-25 Traditional Mutation-Based Fuzzer
For Rust, we adopt rustc [^8] https://www.rust-lang.org/ (the official Rust compiler) and gccrs [^9] https://github.com/Rust-GCC/gccrs/ (the GNU front-end for Rust). For C++, we use GCC [^10] https://gcc.gnu.org/ and Clang, [^11] https://clang.org/, which are the two most widely deployed C++ compilers in practice. For the Rust compilers, we adopt the latest release versions, i.e., rustc 1.88 and gccrs-a1a56c6. For the C++ compiler, we follow the existing study (Ou et al., 2024; Wang et al., 2025b), adopting the most recent releases in 2024 (i.e., GCC 14.1.0 and Clang 18.1.0). As test oracles, we apply crash, hang, and differential testing for compilers of the same language (Gao et al., 2025; Zhang et al., 2017). Among them, crash and hang are used to detect compiler failures. The former indicates a crash or assertion failure, while the latter indicates an infinite loop or deadlock. Differential testing is used to detect inconsistent behaviors between different compilers, which is a common practice for testing C/C++ compilers that share the same language standard but may have different implementations. For the baseline fuzzers, we use their default configurations from their official repositories. For approaches that require seed programs, we use the same initial seed corpus as Mut4All to ensure a fair comparison.
@@ 4.2.1. Baselines
To thoroughly evaluate the effectiveness of Mut4All, we compare it against both traditional compiler fuzzers and recent LLM-based approaches. In particular, to investigate how our mutator-driven pipeline compares to direct LLM-based mutated program generation, we include the latest LLM-based fuzzers that modify program source code directly—namely, Clozemaster (Gao et al., 2025) for Rust and Fuzz4All (Xia et al., 2024) for C++. In total, we evaluate three fuzzers for Rust and six for C++, as shown in Table 1.
Traditional Fuzzers. For Rust, we adopt generation-based fuzzers such as RustSmith (Sharma et al., 2023) and Rustlantis (Wang and Jung, 2024). For C++, we include CSmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023), which follow a similar generation-based paradigm by randomly expanding language grammars and production rules to produce valid programs. We include GrayC (Even-Mendoza et al., 2023), a representative mutation-based C++ fuzzer equipped with manually designed, general-purpose mutators. Additionally, we adopted the most recent C++ fuzzer, TyMut (Wang et al., 2025b), which focuses on the type system of C++ compilers.
LLM-based Fuzzers. We further compare Mut4All with recent LLM-based compiler fuzzers that follow different design philosophies: we let the LLMs design and implement mutators, which are then applied to the seed programs, but they let the LLMs directly edit the source code to generate mutated programs. For Rust, Clozemaster (Gao et al., 2025) fine-tunes an LLM to complete masked code segments, enabling direct in-place edits to source programs. For C++, Fuzz4All (Xia et al., 2024) utilizes LLMs to rewrite seed programs guided by language standards and official API references. In contrast, MetaMut (Ou et al., 2024) uses an LLM to systematically synthesize and combine simple transformation rules into new mutators, falling into the same category as Mut4All but with limited mutator diversity.
@@ 4.3. RQ1: Comparison with Existing Fuzzers
To evaluate Mut4All against existing approaches, we run each fuzzer with a 24-hour time budget, which is a common setting in prior work (Even-Mendoza et al., 2023; Xia et al., 2024; Ou et al., 2024; Yang et al., 2024a; Wang et al., 2025b). Also, each fuzzer is executed independently for five trials. Following existing studies (Ou et al., 2024; Wang et al., 2025b), we report the average of the line coverage within the multiple rounds, and collect the union of all recovered crashes during all rounds. For fare comparison, Fuzz4All and MetaMut both use the same backend LLM, GPT-4o, as Mut4All, while Clozemaster is equipped with their original fine-tuned model.
@@ 4.3.1. Coverage
Figure 6. The average line coverage of compilers within 5 rounds of 24h-fuzzing.
Figure 6 presents the average coverage trends of all evaluated fuzzers. For Rust (Figure 6(a)), Mut4All achieves the highest line coverage on rustc, outperforming RustSmith, Clozemaster, and Rustlantis. For C++ compilers, Mut4All achieves the highest code coverage on both GCC (Figure 6(b)) and Clang (Figure 6(c)), followed by another LLM-based C++ compiler fuzzer, MetaMut. Mut4All not only achieves the highest final coverage but also demonstrates a much faster convergence rate within the first two hours, reaching near-saturation significantly earlier than all baselines. The smoother curves further indicate more stable exploration behavior across runs. Moreover, the consistent top performance across both Rust and C++ compilers highlights the cross-language generality of our automated, multi-agent mutator synthesis framework. Overall, these results confirm that Mut4All not only matches but surpasses state-of-the-art fuzzers in improving coverage efficiency and stability for complex compilers.
4.3.2. Unique Crashes
Following the practice of existing studies (Ou et al., 2024; Wang et al., 2025b), a crash is considered unique if its top two stack trace frames—excluding compiler helper functions—differ from those of all other crashes. We aggregate all unique crashes across the five runs for each fuzzer, as shown in Figure 7.
Mut4All consistently discovers the highest number of unique crashes for both Rust and C++, significantly surpassing all baselines. For Rust (Figure 7(a)), Mut4All uncovers 60 unique crashes, whereas Clozemaster and Rustlantis report only 5 and 0, respectively. Notably, only 6 crashes overlap between Mut4All and Clozemaster, indicating that the vast majority of bugs found by Mut4All are exclusive. For C++, Mut4All detects 45 unique crashes on both compilers. The overlap of TyMut, MetaMut, Fuzz4All, and Mut4All is minimal, only 4 crashes are shared among Mut4All and each of these fuzzers. Particularly, Mut4All uniquely recovers 13 crashes that are not covered by all other fuzzers, demonstrating it is highly complementary to other approaches.
(a) The Rust compiler rustc.
(b) The C++ compilers GCC and Clang.
Figure 7. Venn diagram of discovered unique crashes.
Answer to RQ1: Mut4All achieves the best coverage on Rust and both C++ compilers. For the unique crash, Mut4All recovers 60/13 crashes in Rust/C++ compilers that are not covered by the others, indicating it is highly complementary to existing fuzzers.
@@ 4.4. RQ2: Bug-Hunting Effectiveness
We have synthesized 319 and 403 valid mutators for Rust and C++, respectively. Using these mutators, we conducted a long-term fuzzing campaign for each compiler (rustc and gccrs for Rust, and GCC and Clang for C++).
@@ 4.4.1. The Status of Our Reported Bugs
Based on the bug-hunting experiment, we reported a total of 96 bugs, as summarized in Table 4.4.1. Note that a unique crash in the previous RQ does not necessarily indicate a newly discovered bug, as some crashes may correspond to issues that have already been fixed in recent compiler versions. Finally, we identified 62 bugs in Rust compilers (44 in rustc and 18 in gccrs) and 34 in C++ compilers (21 in GCC and 13 in Clang). Among these, 58 bugs (i.e., 60%) have been confirmed by compiler maintainers as previously unknown bugs: 32 in rustc, 9 in gccrs, 7 in GCC, and 10 in Clang. Notably, 22 of the confirmed bugs have already been fixed by the respective compiler developers. In addition, 15 of our reports were classified as duplicates of existing bugs. For rustc, 2 bugs involve internal compiler features not intended for public use, and were marked as Won’t Fix. In total, there are 21 bugs that are still unconfirmed by the developer for all compilers.
@@
Table 2. Status of the reported bugs by Mut4All.
@@
Compiler rustc gccrs GCC Clang
@@
Confirmed as New 32 9 7 10 58
Fixed 12 7 2 1 22
Duplicate 10 0 2 3 15
@@ 4.4.2. Bug Examples
We further present several more bugs to show the effectiveness of Mut4All in uncovering subtle, hard-to-construct compiler issues beyond prior fuzzers.
@@
4// ‘Trait<V, D>‘ -> ‘Trait<NonExistentTrait, NonExistentTrait>‘
5impl<V, D> Trait<NonExistentTrait, NonExistentTrait> for A
6where
7 A: Trait<D, V>,
8 B: Trait<V, D>,
9{}
10// ‘Trait<V, D>‘ -> ‘Trait<NonExistentTrait, NonExistentTrait>‘
11impl<V, D> Trait<NonExistentTrait, NonExistentTrait> for B where A: Trait<V, D> {}
12fn impls_trait<T: Trait<V, D>, V, D>() {}
13fn main() { impls_trait::<B, _, _>(); }
Listing 7: The new bug Rust-142863 uncovered by Mut4All.
Listing 7 shows a Rust program in which valid type parameters in trait implementations are replaced with a deliberately undefined type (i.e., NonExistTrait). This seemingly illogical mutation causes rustc to enter an infinite compilation loop due to unresolved recursive bounds. Manually crafting such a case would be highly unintuitive, as developers typically avoid inserting undefined types into generic contexts due to their low likelihood of type-checking. However, through automated mutator design and synthesis, Mut4All identifies this as a fruitful transformation, revealing a deep corner case in the compiler’s trait resolution logic and its mishandling of cyclic dependencies introduced by invalid substitutions.
1void copy5CharsInto(char *buf)
2 __attribute__((enable_if(
3 // The original code: ‘__builtin_object_size(buf, 1) != -1 &&‘
4 __builtin_object_size(buf, 1024LL) != -1024LL &&
5 // The original code: ‘__builtin_object_size(buf, 1) > 5,‘
6 __builtin_object_size(buf, 1024LL) > 1024LL,
7 "")));
Listing 8: The new bug LLVM-147009 uncovered by Mut4All.
@@ 8 B: Trait<V, D>,
Listing 9: The new bug LLVM-147170 uncovered by Mut4All.
Listing 8 presents a C++ program in which integer literals used in built-in size checks are replaced with a large edge-case value (1024LL). This subtle transformation triggers a long-standing crash in Clang that had silently persisted since version 9.0, remaining undetected for nearly six years. Although the modified conditions may appear redundant or even nonsensical from a developer’s perspective—since the second argument to the built-in checks is typically expected to be between 0 and 3—this mutation exposes latent bugs in how Clang handles unexpected constant values during attribute evaluation. This example illustrates how Mut4All can move beyond human intuitions and systematically explore unconventional input spaces to uncover deep-seated, previously unreachable compiler defects.
Listing 9 shows a C++ program combining advanced features such as template instantiation and constexpr evaluation. A mutator generated by Mut4All modifies the valid function declaration fooT() into one with an invalid parameter list: fooT(C auto,). While this clearly introduces a syntax error, it unexpectedly causes the Clang compiler to crash rather than emit a diagnostic. This subtle transformation exposed a long-standing bug in Clang’s template and attribute evaluation pipeline that had persisted silently since version 9.0. Such non-obvious mutations—where ill-formed syntax is injected into semantically sensitive constructs—are difficult for traditional mutation-based fuzzers, including MetaMut, to uncover. Moreover, the mutated code appears nonsensical to human developers, who would rarely write deliberately malformed parameter lists in constexpr templates. Mut4All ’s ability to systematically explore these unintuitive transformations highlights its strength in revealing deeply buried compiler defects.
@@
Answer to RQ2: Mut4All discovered 96 bugs across four production compilers from Rust and C++, 58 of which have been confirmed as previously unknown, demonstrating its effectiveness in hunting real compiler bugs. Among these confirmed bugs, 22 of them have already been fixed by the developers.
4.5. RQ3: The Costs for Synthesizing Mutators
We process a total of 1,000 historical compiler bug reports, including 500 for Rust and 500 for C++, through the full multi-agent pipeline of Mut4All. This process ultimately produces 319 valid mutators for Rust and 403 for C++.
To assess the cost-efficiency of Mut4All, we monitor the number of prompts (LLM inputs) and completions (LLM outputs) generated by each agent throughout the synthesis process. As summarized in Table 3, the complete pipeline consumes approximately 4.77M and 8.98M LLM input tokens for Rust and C++, respectively, resulting in 1.33M and 1.15M generated responses.
To further understand how individual agents contribute to the overall synthesis efficiency and how each consumes tokens relative to its productive output, we analyze their respective output yields, defined as the ratio of output tokens to input tokens. As shown in Table 3, the three agents display distinct efficiency characteristics. For Rust, the Mutator Implementation Synthesis agent shows the lowest (accounting for 17%), while for C++ the Mutator Invention agent cost the lowest (accounting for 14%). The difference is cause by the language features and the complexity of the implementation of mutators. For both language, the Mutator Refinement agent demonstrates the highest token costs, which account for 47% for Rust and 57% for C++, respectively. This is expected as the LLMs have to iteratively correct the invalid mutators, which may involve syntax errors, type mismatches, or even incorrect code logic. For all agents, they cost significant more input tokens than output tokens, indicating that an effective prompt engineering approach is necessary to reduce the overall costs.
We further normalize the cost by the number of valid mutators produced. On average, synthesizing a single mutator requires 12,208 LLM tokens for Rust and 20,264 tokens for C++. When considering only valid mutators—and treating the cost of invalid ones as unavoidable overhead—Mut4All consumes 19,135 tokens per Rust mutator and 25,142 tokens per C++ mutator. These figures reflect the total prompt–completion tokens required for fully automated mutator generation and remain practical in modern development pipelines, especially given the reusability of high-quality mutators across multiple fuzzing campaigns. Based on GPT-4o pricing in 2025, the estimated monetary cost for producing one valid mutator is approximately \$0.074 for Rust and \$0.080 for C++. These results indicate that Mut4All provides a cost-effective way to generate a large number of useful mutators efficiently, thereby facilitating early detection of compiler bugs.
Table 3. The costs of each agent of Mut4All.
Language Rust C++
Token Type Input Output Sum Input Output Sum
Mut. Invent. 2,034,827 150,476 2,185,303 1,276,980 147,223 1,424,203
Mut. Impl. Synth. 749,182 290,288 1,039,470 2,694,638 284,096 2,978,734
Mut. Ref. 1,991,245 888,007 2,879,252 5,005,799 723,354 5,729,153
Total 4,775,254 1,328,771 6,104,025 8,977,417 1,154,673 10,132,090
Total / All Mutators # 9,551 2,658 12,208 17,955 2,309 20,264
Total / Valid Mutators # 14,969 4,165 19,135 22,276 2,865 25,142
Answer to RQ3: Mut4All incurs low generation costs: on average, synthesizing a valid mutator requires 14,969/22,276 input tokens and 4,165/2,865 output tokens for synthesizing Rust and C++ mutators, respectively. With GPT-4o pricing in 2025, this corresponds to only \$0.074 per Rust mutator and \$0.080 per C++ mutator.
4.6. RQ4: Mutator Synthesis Effectiveness